OpenStudy (anonymous):

The Law of Sines can only be used to solve problems if the triangle is a right triangle. True False

OpenStudy (anonymous):

The Law of Sines can be used for all triangles

OpenStudy (anonymous):

wait i thought it only applies to 90 degrees triangles

OpenStudy (anonymous):

for every triangle like this law of sins state that  \[\frac{a}{\sin A}=\frac{b}{\sin B}=\frac{c}{\sin C}\]